The driver-side wiper blade is 26 inches long and the passenger-side wiper blade is 16 inches long.
If the surface area of the windshield is 1250 square inches and both blades rotate at an angle of 100°, what percentage of the entire windshield is cleaned by the wiper blades?

Answer by Alan3354(69443)

The driver-side wiper blade is 26 inches long and the passenger-side wiper blade is 16 inches long.
If the surface area of the windshield is 1250 square inches and both blades rotate at an angle of 100°, what percentage of the entire windshield is cleaned by the wiper blades?
Do they overlap? Most do.
What's the distance between centers?
------------
If they don't overlap:
---
Area of the circles pi*r^2 = pi*(26^2 + 16^2) = 932pi sq inches.
The wiper's area = 932pi*100/360 = 2330pi/9 sq inches.
---
% = (100*2330*pi/9)/1250